Parth (parthkohli):

Quick question. Prove that an arbitrary set \(\mathbf{A_n}\) where \(n\) is the number of elements has a power-set with \(2^n\) elements.

Parth (parthkohli):

Induction, but how?

Parth (parthkohli):

Hmm, let's involve \(n = 1 \) to start with.\[\mathcal{P}\{\emptyset\} =\{ \emptyset \} \]This has one element. True for \(n =1\).

Parth (parthkohli):

I still don't understand how to go about with assuming that this is true \(n = k\) and \(k + 1\) or not.

Parth (parthkohli):

@amistre64 Your assistance is certainly needed, sire.

OpenStudy (anonymous):

Sorry for Not -my origin but anyway -wkp

OpenStudy (anonymous):

As you see the argument is as follows: For n+1 the power set will be twice as numerous because - there will be 2 copies of each PREVOUS subset. One "copy" not including a new (n+1)-th member and one copy that does include the new member

OpenStudy (anonymous):

Mr asker @ParthKohli ?

OpenStudy (anonymous):

I get it - the Olympians reside in the olympus, and me - well not an olympian obviously.

OpenStudy (amistre64):

by proving it for some specific "n", we are only proving it for one value; we assume its true for any k, which allows k to be equal to n at some point. proving that it is true for k+1 allows us to deduce that since k=n at some point, then k+1 = n+1 as well

OpenStudy (anonymous):

Thanks @ganeshie8 , aren't you afraid to give a medal to अछूत ?

Parth (parthkohli):

Well, is proving by using combinations a good idea?

ganeshie8 (ganeshie8):

lol i liked the logic given :) 2^n + 2^n thingy... :)

OpenStudy (amistre64):

combinations is proved by induction if i recall correctly:)

Parth (parthkohli):

Does there even exist such a proof? For example, if a set has \(4\) elements, then its powerset will have\[\sum_{i =0}^4 {^4C_i}\]

Parth (parthkohli):

elements.

Parth (parthkohli):

well, i = 1 is the lower limit.

OpenStudy (anonymous):

The elements of power set 2^S can be put in 1:1 correspondence to characteristic functions from S to {Yes include, Not Include} = {1, 0} . Obviously the number of such choices is \[2^{|S|}\]

OpenStudy (amistre64):

basis: \[P\{A_0\} = \{\ \} = 2^0\] assume:\[P\{A_k\} = \{\{\ \},\{1\},\{2\},\{3\},...,{} \} = 2^k\] lol, showing a generic setup for that does look a bit tricky

OpenStudy (amistre64):

{},{1},{2},{3},...,{k}, {1,2},{1,3},{1,4},...,{1,k}, {2,3},{2,4},{2,5},...,{2,k}, pfft .....

OpenStudy (amistre64):

Would you have to reinvent how to define a power set?